# Prompt Injection Vulnerability of Consensus Generating Applications in Digital Democracy

### v3

@@ 6. AMBS, University of Manchester
Large Language Models (LLMs) are gaining traction as a method to generate consensus statements and aggregate preferences in digital democracy experiments. Yet, LLMs could introduce critical vulnerabilities in these systems. Here, we examine the vulnerability and robustness of off‑the‑shelf consensus‑generating LLMs to prompt‑injection attacks, in which texts are injected to amplify particular viewpoints, erase certain opinions, or divert consensus toward unrelated or irrelevant topics. We construct attack‑free and adversarial variants of prompts containing public policy questions and opinion texts, classify opinion and consensus valences with a fine‑tuned BERT model, and estimate Attack Success Rates (ASR) from 3×3 confusion matrices conditional on matching human majorities. Across topics, default LLaMA 3.1 8B Instruct, GPT‑4.1 Nano, and Apertus 8B exhibit widespread vulnerability, with especially high ASR for economically and socially conservative parties and for rational, instruction‑like rhetorical strategies. A robustness pipeline combining GPT‑OSS‑SafeGuard injection detection, structured opinion representations, and GSPO‑based reinforcement learning reduces ASR to near zero across parties and policy clusters when restricting attention to non‑ambiguous consensus outcomes. These findings advance our understanding of both the vulnerabilities and the potential defenses of consensus‑generating LLMs in digital democracy applications.
@@ 1 Introduction
Because of their ability to process, classify, and rank vast amounts of textual information, Large Language Models (LLMs) have become a popular tool among researchers exploring the use of AI in digital democracy (DD) applications (Tessler et al. 2024; Ash, Galletta, and Opocher 2025; Gudiño, Grandi, and Hidalgo 2024; Li et al. 2025; Konya et al. 2025; Majumdar, Elkind, and Pournaras 2024; Small et al. 2023). In these systems, LLMs are used to summarize arguments, predict preferences, or converse with citizens to help them explore policy options. Yet, while LLMs can facilitate many of these tasks, its use comes with important limitations (Helbing and Sánchez-Vaquerizo 2023; García-Marzá and Calvo 2025; Novelli et al. 2025).
LLMs have vulnerabilities that attackers are likely to target. We can divide these attacks roughly into two categories. Attacks involving user level access, such as prompt-injection and information extraction attacks (Qi et al. 2025; Mattern et al. 2023), and attacks requiring higher levels of access, such as data poisoning or application parameter attacks (Zhang et al. 2024,; Berdoz et al. 2025).
@@ 2 Related work
During the last few years, prompt injection attacks have become a growing concern in the use of LLMs. There is evidence of prompt injection attacks where academics by hide statements in papers with the goal of manipulating LLM generated reviews (Gibney 2025) and there are concerns about prompt-injections being hidden in the data retrieved by RAG systems (Qi et al. 2025). Unfortunately, because of the flexibility of language, it is difficult to enumerate all possible forms of prompt-injection attacks.
@@ 3 Prompt-Injection Attacks in LLM-based Democratic Deliberation
Human/Machine readable is rather self-explanatory, as it separates prompts that can be read by humans (Chen et al., 2025b) and those that are designed for machines, which look more like code (Pasquini, Strohmeier, and Troncoso, 2024). In this paper, we focus on human-readable prompt injections.
@@ 4.2 Testing Vulnerability to Prompt Injections
To test whether consensus generation is vulnerable to prompt injection attacks, we begin by drawing a random sample of prompts along with their respective opinion texts. These prompts are linked to 62 (out of 301) public policy questions. We classified each opinion text into one of three valences (”Agree” with the question–green–, ”Disagree” –red–, or ”Ambiguous”–yellow–, in Figure 2) using a BERT classifier fine-tuned on GPT4o labels (F1=0.98, see Appendix B of the supplementary material for details). We use the BERT classifier instead of using the GPT4o labels directly, for cost considerations.
@@
Then, we pair each of the consensus statements generated using attack-free data with their corresponding prompt injected consensus to produce a 3×3 confusion matrix. Here, rows correspond to the valence of the consensus statement without prompt injections (agree, disagree, or ambiguous consensus) and columns correspond to the same valences under prompt injections.
@@
To ensure that changes in consensus outcomes reflect the effect of prompt injections rather than LLM performance, we restrict our analysis to pairs where the original prompt yields an LLM-consensus whose valence (agree, disagree, ambiguous) matches the majority view of participants under a majority-with-abstentions rule (see Appendix C for precise definition of this rule and Appendix D for default LLM performance). This rule covers edge cases, such as evenly split groups (2–2) or cases where the majority itself expresses ambiguity. Applying this filter yields a dataset of 35,521 consensus pairs with overlapping categories: 17,403 with support attacks, 18,049 with criticism attacks, 17,761 with completion attacks and 17,761 with ignore attacks.
Finally, we compute our evaluation metric: the Attack Success Rate (ASR) across each ignore/completion, framing and rhetorical strategy. ASR measures the proportion of consensus pairs in which a single prompt injection shifts the LLM’s valence away from what it would have produced in the absence of attacks. While attacks are designed to be directional—either supporting or criticizing the policy embedded in the question—this metric also captures unintended shifts, where the valence changes in ways not anticipated by the attack. In the 3x3 confusion matrix illustrated in Figure 2, the ASR is the share of consensus pairs that fall outside the diagonal.
@@
In the presentation of our results, we aggregate ASR values across three key dimensions (see Appendix E for details). First, we analyze clusters of public policies using sentence embeddings to assess whether certain types of public policies are systematically more vulnerable to prompt injections. Second, we relate each public policy question to political parties’ stated positions by extracting relevant paragraphs from their 2024 manifestos using DeepSeek-OCR (Wei, Sun, and Li 2025) and RAG, and compute party-specific ASRs depending on whether a party supports or opposes the policy. Although parties played no role in the deliberation process, linking consensus shifts to manifesto positions shows which political actors would benefit from successful attacks. For instance, if an attack successfully shifts consensus from ’Disagree’ to ’Agree’ on a healthcare policy question, any political party whose manifesto supports that policy would benefit from the attack. Third, we aggregate ASR by rhetorical strategy to identify the most effective persuasive techniques at shifting consensus outcomes.
@@ 4.3 Testing Robustness against Prompt Injections
As a first step, we identify whether each opinion and injection text linked to the training prompts contains a prompt injection, using GPT-OSS-SafeGuard (Agarwal et al. 2025). All positive detections are replaced with an “Opinion Deleted” tag in the corresponding prompt. This procedure allows us to leverage recent advances in AI cybersecurity benchmarks while enabling rapid localization of potential injections within prompts—a feature of particular importance for deliberative democracy systems with many participants. Using this approach, 99.31% of injection texts and 0% of opinion texts are flagged as attacks, outperforming Qwen3Guard and Syntactic Dependency Parsing (see Appendix F).
@@
Figure 4: Vulnerability (panel (a)) and robustness (panel (b)) to prompt‑injection attacks for LLaMA 3.1 8B Instruct, expressed as Attack Success Rates (ASR) by public policy question cluster (x‑axis) and UK political party (y‑axis). Panel (a) reports baseline vulnerability, with ASR computed from the disagree, ambiguous, and agree cells of the underlying confusion matrices, and all elements shown in red. Panel (b) displays clustered bar charts, where red bars correspond to ASR without robustness (using only the agree and disagree categories) and blue bars correspond to ASR after applying the robustness methodology with the same categories. In both panels, the top bar chart displays the bootstrap mean ASR for each public policy question cluster, the bar chart on the right displays the bootstrap mean ASR for each political party, and the bottom‑left heatmap reports bootstrap mean ASR for every party–cluster pair, with darker shades indicating higher attack success and black cells indicating insufficient data. All ASR values and their 95% confidence intervals are estimated via bootstrapping with 5,000 iterations. Appendix H reports analogous results for GPT 4.1 Nano and Apertus 8B.
Next, we merge the prompts of the training set containing these structured representations and the “Opinion Deleted” tags into a unified dataset. Because default LLMs perform poorly on the arithmetic and aggregation operations required to implement our majority rule, this unified dataset is used to align models through Group Sequence Policy Optimization (GSPO) (Zheng et al. 2025), following the hyperparameters and reward functions of Pappone et al.. GSPO fine-tunes LLMs via reinforcement learning, explicitly rewarding group-level sequence generations that internally satisfy the majority rule (see Appendix G for details). The resulting system functions as an opinion aggregation and filtering pipeline that prioritizes intended deliberative outputs—such as majority-consistent summaries—even in the presence of adversarial perturbations. While future methods may offer stronger guarantees (Debenedetti et al. 2025; Cao et al. 2025), our goal here is to assess the feasibility of building resilient deliberative-democracy systems with the most effective techniques currently available.
Finally, we evaluate the aligned aggregation pipeline using the same procedures and result format described in the vulnerability analysis. The same structured transformation is applied to the prompts of the test set, ensuring methodological consistency between training and evaluation. A lower Attack Success Rate (ASR) after fine-tuning indicates that the system more consistently preserves its intended deliberative behavior under prompt-injection attacks.
@@ 5 Results
5.1 Vulnerability to Prompt-Injections
Figure 4 (panel (a)) maps the vulnerability of consensus generation in the default LLaMA 3.1 8B Instruct model to prompt-injection attacks across policy domains and political representations. Attack success rates (ASR) are computed from the disagree, ambiguous, and agree cells of the underlying confusion matrices and are decomposed along three dimensions: clusters of public policy questions (top bar chart), UK political parties (right bar chart), and their intersection (heatmap).
At the level of policy domains, susceptibility to prompt injections is remarkably homogeneous. Most policy question clusters exhibit ASR values between roughly 26% and 39%, indicating that consensus generation is broadly vulnerable across topics, rather than being driven by a small set of particularly fragile domains. In isolation, no policy area appears structurally immune to manipulation.
The picture changes when attention shifts to political representations. Attack effectiveness is highly uneven across parties, with unionist and socially conservative parties—such as Reform UK, the Conservative Party, and Ulster/Democratic/Traditional unionists—showing substantially higher ASR values, ranging from about 41% to nearly 80%. In contrast, the remaining parties exhibit markedly lower vulnerability, with ASR generally below 27%. This asymmetry suggests that political framing and ideological structure play a stronger role in mediating vulnerability than policy content alone.
The joint distribution of policy domains and political parties further sharpens this pattern. The heatmap reveals that prompt-injection effectiveness is concentrated in parties with conservative economic and social orientations, particularly in discussions on green policy—e.g. Conservatives and Ulster/Democratic Unionists, where ASR frequently exceeds 50%—and on public spending and taxation—e.g., Ulster/Democratic Unionist—, where values often surpass 30%. By contrast, green, nationalist (such as the SNP and Plaid Cymru), and progressive parties show relatively higher vulnerability in debates concerning democratic rights and social control, with ASR values around 40%, and more moderate susceptibility in green policy discussions, with ASR values between 10% and 30%.
Taken together, these results suggest that prompt-injection attacks exploit structured asymmetries in political representation and, to a lesser degree, specific policy topics.
Although the detailed patterns differ across policy clusters and within the heatmap, this asymmetry across political parties is replicated in additional experiments using GPT-4.1 Nano and Apertus 8B (see Appendix H). With GPT-4.1 Nano, TUV, Reform UK, and the DUP register the highest attack effectiveness, with ASR exceeding 40%. Under Apertus 8B, conservative and unionist parties exhibit ASR values between 30% and 50%, while all others remain below 25%.
In Appendix I, we break down attack effectiveness along the rhetorical-strategy axis of our taxonomy. Explicit instructions and rational-sounding arguments—captured by the strategies Imperative Order and Impossibility of Agreement—prove consistently more effective, yielding ASR values between 35% and 52%. In contrast, emotional language and fabricated statistics lag behind across all the LLMs we evaluate, with ASR values in the 25%–32% range.
To assess whether our vulnerability findings scale to larger and more heterogeneous settings, we replicated in Appendix J the analysis using 26,502 candidate statements from SmartVote, a platform that surveys political candidates during Switzerland’s parliamentary elections, covering 374 policy questions (Stammbach et al. 2024). For each combination of party, language (German, French, or Italian), and policy question, we generated within-party consensus statements by aggregating 5–40 opinion texts, yielding 48,192 test prompts varying in framing, ignore/completion structure, and rhetorical strategy. ASR values remain above zero even with more than 30 contributors, confirming that single attacks persist in both larger and linguistically heterogeneous groups.
@@ 5.2 Robustness Against Prompt Attacks
Figure 4 (panel (b)) decomposes the effectiveness of defense techniques against prompt-injection attacks in LLaMA 3.1 8B Instruct across policy domains and political representations. To focus on clearly directional outcomes, Attack Success Rates (ASR) are computed from a reduced 2×2 confusion matrix over Agree and Disagree consensus valences, excluding the Ambiguous category. As observed in Appendix K, incorporating Ambiguous outcomes into the full 3×3 specification leads to much higher ASR estimates, reflecting the fact that small perturbations can easily flip intrinsically uncertain cases; for this reason, we concentrate our main robustness analysis on non-ambiguous consensus outcomes. Results are presented along three dimensions: clusters of public policy questions (top bar chart), UK political parties (right bar chart), and their intersection (heatmap).
The heatmap and the clustered bar charts show that once our defense technique is applied, ASR values fall close to 0% across nearly all political parties and policy clusters. The only outliers are People Before Profit and Sinn Féin in debates tied to green policy and sustainable energy—yet even there ASR values remain below 5%. When dissecting ASR across the stages of our methodology (see Appendix K), we find that filtering and structured opinion representations bring ASR down to the 1.8%–11.8% range, and GSPO pushes those values even further, landing between 0% and 1.9%.
Within the broader landscape of defense strategies (Chen et al. 2025a, b), we benchmark our approach against two baselines: a combination of Direct Preference Optimization (DPO) (Rafailov et al. 2023) and GSPO (Appendix L), and Deliberative Alignment (Guan et al. 2024) (Appendix M). Both deliver solid but consistently lower performance, despite operating directly on the original texts rather than on transformed representations. DPO and GSPO reduce attack effectiveness to 3.2%–23% across political parties and 2.2%–44% across public-policy clusters. Deliberative Alignment brings these values down to 5.3%–21.1% across political parties and 1.9%–34.4% across public-policy clusters.
An over-refusal rate analysis (see Appendix N), comparing consensus statement valences with and without defense techniques under absence of prompt injection attacks, shows that our approach yields a very low overrefusal risk, with ASR values ranging between 1.2% and 3.7% after applying GSPO.

### v4

@@ 6. AMBS, University of Manchester
Large Language Models (LLMs) are gaining traction as a method to generate consensus statements and aggregate preferences in digital democracy experiments. Yet, LLMs could introduce critical vulnerabilities in these systems. Here, we examine the vulnerability and robustness of off‑the‑shelf consensus‑generating LLMs to prompt‑injection attacks, in which texts are injected to amplify particular viewpoints, erase certain opinions, or divert consensus toward unrelated or irrelevant topics. We construct attack‑free and adversarial variants of prompts containing public policy questions and opinion texts, classify opinion and consensus valences with a fine‑tuned BERT model, and estimate LLM-human majority agreement rates. Across topics, default LLaMA 3.1 8B Instruct, GPT‑4.1 Nano, and Apertus 8B exhibit widespread vulnerability, specially when disagreement and disagreement are finely balanced, for attacks that shift consensus toward positions aligned with GB-unionist conservative manifestos relative to pro-independence left manifestos, and for rational, instruction-like rhetorical strategies. A robustness pipeline combining GPT‑OSS‑SafeGuard injection detection, structured opinion representations, and GSPO‑based reinforcement learning substantially reduces directional failures whenever the underlying consensus has a clear positive or negative valence. These findings advance our understanding of both the vulnerabilities and the potential defenses of consensus‑generating LLMs in digital democracy applications.
@@ 1 Introduction
Because of their ability to process, classify, and rank vast amounts of textual information, Large Language Models (LLMs) have become a popular tool among researchers exploring the use of AI in digital democracy (DD) applications (Tessler et al. 2024; Ash, Galletta, and Opocher 2025; Gudiño, Grandi, and Hidalgo 2024; Li et al. 2025; Konya et al. 2025; Majumdar, Elkind, and Pournaras 2026; Small et al. 2023). In these systems, LLMs are used to summarize arguments, predict preferences, or converse with citizens to help them explore policy options. Yet, while LLMs can facilitate many of these tasks, its use comes with important limitations (Helbing and Sánchez-Vaquerizo 2023; García-Marzá and Calvo 2025; Novelli et al. 2025).
LLMs have vulnerabilities that attackers are likely to target. We can divide these attacks roughly into two categories. Attacks involving user level access, such as prompt-injection and information extraction attacks (Mattern et al. 2023), and attacks requiring higher levels of access, such as data poisoning or application parameter attacks (Zhang et al. 2024,; Berdoz et al. 2025).
@@ 2 Related work
During the last few years, prompt injection attacks have become a growing concern in the use of LLMs. There is evidence of prompt injection attacks where academics by hide statements in papers with the goal of manipulating LLM generated reviews (Gibney 2025) and there are concerns about prompt-injections being hidden in the data retrieved by RAG systems (Clop and Teglia 2024). Unfortunately, because of the flexibility of language, it is difficult to enumerate all possible forms of prompt-injection attacks.
@@ 3 Prompt-Injection Attacks in LLM-based Democratic Deliberation
Human/Machine readable is rather self-explanatory, as it separates prompts that can be read by humans (Chen et al., 2025) and those that are designed for machines, which look more like code (Pasquini, Strohmeier, and Troncoso, 2024). In this paper, we focus on human-readable prompt injections.
@@ 4.2 Testing Vulnerability to Prompt Injections
To test whether consensus generation is vulnerable to prompt injection attacks, we begin by drawing a random sample of prompts along with their respective opinion texts. These prompts are linked to 62 (out of 301) public policy questions. We classify each opinion text into one of three valences (”Disagree” –red–, ”Ambiguous” –yellow–, or ”Agree” with the question –green– in Figure 2) using a BERT classifier fine-tuned on GPT4o labels (F1=0.98, see Appendix B of the supplementary material for details), which we use in place of GPT4o labels for cost reasons. We further map these valences to numerical values (1-1, 0, and +1+1, respectively) and aggregate them at the prompt level by averaging across opinions. The resulting net position ranging between 1-1 and 11—for instance, 0.4-0.4 in Figure 2— captures both the direction and intensity of collective support, while remaining comparable across prompts with different numbers of participants.
@@
Then, we pair each consensus statement generated from the original prompt with the consensus generated from its injected counterpart, and group these pairs by the prompt’s net position.
@@
To ensure that changes in consensus outcomes reflect the effect of prompt injections rather than LLM performance, we restrict our analysis to pairs where the original prompt yields an LLM-consensus whose valence (disagree, ambiguous, agree) matches the net position, depending if this value is negative (disagree), zero (ambiguous) or positive (agree) (see Appendix C for default performance of LLMs against this calculation).
Finally, we compute as our evaluation metric the LLM agreement rate (AR). For a given a net position, AR is the fraction of paired prompts for which the consensus remains unchanged after the injection, that is, the fraction of pairs for which the LLM assigns the same consensus statement’s valence to the original prompt and to its injected counterpart. Higher values signal lower vulnerability. This calculation is illustrated in the lower portion of Figure 2, which summarizes how vulnerable LLM-based consensus generation is across different group configurations. The x-axis places groups of pairs according to their net position, ordering groups from those dominated by disagreement to those dominated by agreement. The y-axis reports the AR for each net position, showing how often the LLM preserves its original consensus judgment when one opinion is adversarially modified. While attacks are designed to be directional—either supporting or criticizing the policy embedded in the question—this metric also captures unintended shifts, where the valence changes in ways not anticipated by the attack.
@@
In the presentation of our results, we disaggregate AR values across three key dimensions (see Appendix D for details). First, we analyze clusters of public policies using sentence embeddings to assess whether certain types of public policies are systematically more vulnerable to prompt injections. Second, we relate each public policy question to political parties’ stated positions by extracting relevant paragraphs from their 2024 manifestos using DeepSeek-OCR (Wei, Sun, and Li 2025) and RAG, and compute party-specific agreement rates depending on whether a party supports or opposes the policy. Although parties played no role in the deliberation process, linking consensus shifts to manifesto positions shows which political actors would benefit from successful attacks. For instance, if an attack successfully shifts consensus from ’Disagree’ to ’Agree’ on a healthcare policy question, any political party whose manifesto supports that policy would benefit from the attack. Third, we aggregate agreement rates by rhetorical strategy to identify the most effective persuasive techniques at shifting consensus outcomes.
Figure 4: Vulnerability to prompt-injection attacks measured as the LLM–majority agreement rate ($y$-axis) against net position ($x$-axis, ranging from 1-1 to 11). (a) LLM–majority agreement rate by support/criticism attacks. (b) LLM–majority agreement rate by public policy clusters: institutions and infrastructure (“I”), public spending and taxation (“S”), redistribution and labor rights (“R”), education policy and parental rights (“E”), green policies (“G”), environmental and consumption regulation (“E/C”), democratic rights and social control (“D”). (c) LLM–majority agreement rate by rhetorical strategy: emotional appeals (“E”), false authority (“A”), imperative order (“I”), impossibility of agreement (“NA”), misleading statistics (“S”). (d) LLM–majority agreement rate by LLM. (e)–(f) LLM–majority agreement rate by UK political parties: GB-unionist right (“Cons.”, “Reform UK”) and pro-independence left: Plaid Cymru (“PC”), Green Party (“GP”), Scottish Greens (“SG”), Scottish National Party (“SNP”), Sinn Féin (“SF”), People Before Profit (“PBP”). Results for remaining parties are presented in Appendix I. All agreement rate values and their $95\%$ confidence intervals are estimated via bootstrapping with 5,0005{,}000 iterations.
@@ 4.3 Testing Robustness against Prompt Injections
As a first step, we identify whether each opinion and injection text linked to the training prompts contains a prompt injection, using GPT-OSS-SafeGuard (Agarwal et al. 2025). All positive detections are replaced with an “Opinion Deleted” tag in the corresponding prompt (see Figure 3). This procedure allows us to leverage recent advances in cybersecurity benchmarks while enabling rapid localization of potential injections within prompts—a feature of particular importance for deliberative democracy systems with many participants. Using this approach, 99.31% of injection texts and 0% of opinion texts are flagged as attacks, outperforming Qwen3Guard and Syntactic Dependency Parsing (see Appendix E). Data leakage is unlikely, as no public LLM consensus-generation benchmarks existed prior to its knowledge-cutoff date (June 2024).
@@
Next, we merge the prompts of the training set containing these structured representations and the “Opinion Deleted” tags into a unified dataset. Because default LLMs perform poorly on the arithmetic and aggregation operations required to implement the calculation of the net position, this unified dataset is used to align models through Group Sequence Policy Optimization (GSPO) (Zheng et al. 2025), following the hyperparameters and reward functions of Pappone et al.. GSPO fine-tunes LLMs via reinforcement learning, explicitly rewarding group-level sequence generations that internally satisfy the calculation of the net position (see Appendix F for details). The resulting system functions as an opinion aggregation and filtering pipeline that prioritizes intended deliberative outputs—such as net position-consistent summaries—even in the presence of adversarial perturbations. While future methods may offer stronger guarantees (Debenedetti et al. 2025; Cao et al. 2026), our goal here is to assess the feasibility of building resilient deliberative-democracy systems with the most effective techniques currently available.
Finally, we evaluate the aligned aggregation pipeline using the same procedures and result format described in the vulnerability analysis, as it can be observed in Figure 3. The same structured transformation is applied to the prompts of the test set, ensuring methodological consistency between training and evaluation. A higher LLM Agreement Rate (AR) after fine-tuning indicates that the system more consistently preserves its intended deliberative behavior under prompt-injection attacks.
@@ 5 Results
5.1 Vulnerability to Prompt Injections
Figure 4 summarizes how vulnerable LLM-based consensus generation is to prompt-injection attacks across different group configurations. Along the x-axis, prompts are ordered by their net position, ranging from collective disagreement (1-1) to collective agreement (+1+1). The y-axis reports the LLM–majority agreement rate (AR), measuring how often the LLM preserves its original consensus valence after one opinion is adversarially modified. Lower values signal higher LLM vulnerability to prompt injections.
Figure 5: Robustness to prompt-injection attacks measured as the LLM–majority agreement rate ($y$-axis) against net position ($x$-axis, ranging from 1-1 to 11) after applying the defense pipeline. (a) LLM–majority agreement rate by support/criticism attacks. The no-defense condition is labeled as “ND”. (b) LLM–majority agreement rate by defense layer: no defense (“ND”), GPT-OSSafeGuard combined with structured representations (“G + S”), and the full pipeline adding GSPO fine-tuning (“G + S + GSPO”). (c) LLM–majority agreement rate by methodology: no defense (“ND”), DPO combined with GSPO (“DPO + GSPO”), deliberative alignment (“DA”), and the full pipeline (“G + S + GSPO”). (d)–(f) Distribution of valences of LLM-generated consensus statements after applying the defense pipeline across negative, neutral and positive net positions. All agreement rate values and their $95\%$ confidence intervals are estimated via bootstrapping with 5,0005{,}000 iterations.
Panel (a) contrasts support and criticism attacks against a no-attack baseline (100%) in which the valence of the consensus statement matches the valence of the net position, showing that agreement rates drop sharply at neutral net positions, and, as expected, that criticism (support) attacks are most effective against positive (negative) net positions, that is, when the original consensus leans toward agreement (disagreement).
Panel (b) disaggregates agreement rates by public policy cluster, showing that vulnerability to prompt injections is unevenly distributed across policy domains. Policies related to education, green policies, environmental and consumption regulation exhibit systematically lower agreement rates, especially around neutral net positions. By contrast, clusters such as institutions and infrastructure, public spending and taxation display higher agreement rates across most of the net-position range, indicating more stable consensus formation in these domains.
Panel (c) contrasts rhetorical strategies, revealing large confidence intervals that point to substantial variability in attack effectiveness. Despite this uncertainty, two patterns stand out: imperative-order and impossibility-of-agreement attacks consistently achieve lower agreement rates, especially at positive net positions. This suggests that rational-sounding arguments are more relatively effective than emotionally charged language or fabricated statistics at shifting consensus statements.
Panel (d) shows that these vulnerabilities persist across different LLMs.
Panels (e) and (f) break down agreement rates by party alignment, revealing an asymmetry depending on whether a party’s manifesto supports or opposes public policies: For GB-unionist right parties (UK Conservative and Reform UK, Panel (e)), agreement rates are systematically lower at the extremes of the net-position axis than for pro-independence left parties (green parties—Green Party and Scottish Greens; Welsh and Scottish nationalists—Plaid Cymru and Scottish National Party; Irish republicans—Sinn Féin; and People Before Profit, Panel (f)). Results for other parties are presented in Appendix I.
Overall, these results highlight that prompt-injection attacks exploit structured asymmetries in group composition and framing, producing the largest disruptions when collective preferences are weak rather than strongly aligned.
@@ 5.2 Robustness Against Prompt Attacks
Figure 5 summarizes how effectively the proposed defense pipeline restores LLM-based consensus generation against prompt-injection attacks across different configurations. Along the x-axis, prompts are ordered by their net position, ranging from collective disagreement (1-1) to collective agreement (+1+1). The y-axis reports the LLM–majority agreement rate (AR), measuring how often the LLM preserves its original consensus valence after one opinion is adversarially modified. Higher values signal greater robustness to prompt injections.
Panel (a) contrasts support and criticism attacks against the no-defense baseline, showing that the pipeline neutralizes each attack type where it was most damaging in the vulnerability analysis: AR reaches $100\%$ for support attacks at negative net positions and for criticism attacks at positive net positions. Yet, in the remaining cases AR drops sharply for both attack types, suggesting residual vulnerability.
Panel (b) disaggregates these results by defense layer, showing that the robustness gains in panel (a) are driven primarily by GSPO fine-tuning. GPT-OSSafeGuard combined with structured representations alone yields inconsistent gains, occasionally performing worse than the no-defense baseline near neutral net positions. Adding GSPO substantially raises AR across most of the net-position range, though dips at neutral positions persist.
Panel (c) contrasts the full pipeline against alternative alignment methodologies. Deliberative alignment (Guan et al. 2024) (Appendix H) and DPO (Rafailov et al. 2023) combined with GSPO (Appendix G) both exhibit high variance and repeatedly fall below the no-defense baseline, while the full pipeline maintains consistently higher AR across the net-position range.
Panels (d)–(f) unpack the residual LLM–human majority mismatches by showing, for each net position, how often the pipeline produces consensus statements with disagreement, ambiguous, or agreement valences. For negative or positive net positions ((d) and (f)), the pipeline mostly preserves the original valence (disagreement or agreement), and residual mismatches are almost entirely due to consensus statements becoming ambiguous rather than flipping valence. In contrast, when the valence of the original consensus is ambiguous (net position $\approx 0$, panel (e)), a non-trivial share of consensus statements is pushed into clear agreement or disagreement, indicating that the pipeline has structural difficulty preserving ambiguity even after filtering out prompt-injection attempts. Building LLM-based DD systems that maintain directional protection while also resolving deliberative ambiguity thus emerges as the primary remaining challenge for resilient DD systems.
An over-refusal rate analysis (see Appendix J), measuring how often the defended model changes its valence assessment in the absence of prompt injection attacks, shows that our approach yields a low over-refusal risk. After applying GSPO, over-refusal rates remain consistently below 20% across most net position values when ambiguous cases are included, and drop even further when ambiguity is excluded, staying below approximately 10% for the majority of net positions. In both settings, an isolated peak is observed near 0.5-0.5, reaching approximately 55–80% with ambiguity and around 55% without, which coincides with the most contested consensus positions.